# Universidad de Alcalá, Alcalá de Henares 28805, Spain

### v1

@@ 3 Results
Compared to event 1, event 2 exhibits more complex features, as shown in the right-hand panels of Fig. 2. There are data gaps in the SWA measurements during the pre-event period on 2024 July 22. The proton dynamic spectra in panel (a) also display two distinct tracks of velocity dispersion. Bidirectional streaming characteristics are again apparent in the first-order anisotropy shown in panel (b). In this event, the magnetic field polarity is predominantly inward, so negative first-order anisotropy indicates particles propagating away from the Sun, whereas later positive values correspond to sunward streaming. This interpretation is reflected in panel (c), where the highest intensity of protons initially appears at pitch angles around (as measured by the sunward telescope) but later shifts to around (anti-sunward telescope). Panels (c) and (d) reveal that pitch angle distribution in this event exhibits more pronounced fluctuations than in event 1. The pitch angle distribution of protons in panel (d) shows a feature that is similar to event 1, during the period of 2:00 to 4:00 UT, the highest intensity of protons is measured with pitch angle around but the highest intensity of protons appears at pitch angles around , indicating the bidirectional SEP propagation. Unlike event 1, the bidirectional electron feature measured by EAS is less evident in the pitch angle distribution shown in panel (e). Although the phase space densities near both and are elevated compared to , the peak at is more pronounced in the inward magnetic field configuration. The magnetic field and solar wind data in panel (f)-(i) reveal substantial fluctuations, making it challenging to identify SFRs directly using PyGS. Only a few short-duration and unrelated SFRs are detected. Notably, analysis of the normalized reduced magnetic helicity reveals the presence of a highly helical magnetic structure with a characteristic scale of approximately minutes, spanning from 2024 July 23, around 00:00 to 06:00 UT. Although the precise boundaries are difficult to determine using the wavelet method, the combined results of and in Appendix A.2 indicate that Solar Orbiter crossed a potentially Alfvénic SFR embedded within a turbulent solar wind environment near the onset of the eruption. Unlike event 1, where the SFR was encountered before the eruption, in event 2, the encounter with the SFR is unclear. It seems to occur during the onset phase, which could allow for the observation of bidirectional SEPs. Recent studies have emphasized the prevalence of such SFRs in the slow solar wind (e.g., Chen et al., 2021; Shi et al., 2021), where they exhibit a high degree of Alfvénicity. The possible Alfvénic flux ropes may have a significant impact on SEP transport by affecting both the propagation path and scattering conditions. Unfortunately, we did not find any relevant eruption information associated with this possible SFR in the DONKI database. If the small eruption occurred precisely on the far side of the Sun, it is possible that no evidence would be detectable in coronagraph observations near Earth.
@@
To further understand the transport processes in the two events, we employ VDA to determine the release times and path lengths of the two sources. In the left-hand panels of Fig. 3, the upper panel displays the type III radio burst observed by RPW during event 1. Only a single burst is evident during the onset phase, beginning at approximately 06:41 UT from the Sun, indicated by the vertical line. The lower four panels show proton dynamic spectra from the four HET and EPT telescopes, highlighting the features of velocity dispersion. Two distinct velocity dispersion tracks can be readily identified by visual inspection in both the sunward and anti-sunward telescopes. Notably, the earliest arriving particles detected by the sunward telescopes are scarcely observed by the anti-sunward telescope during the onset. In addition, only a small number of particles are detected in the north and south telescopes at pitch angles near , indicating very limited scattering during the transport process. Compared to the proton measurements, velocity dispersion features are less pronounced in the electron data, likely due to high-level background noise. Therefore, we present the electron measurements by EPT in the Appendix A.1 and do not perform VDA for electrons in this study.
@@
Onset times are determined independently for the sunward and anti-sunward telescopes using the Poisson-CUSUM method (Lucas, 1985). VDA yields an injection time for source 1 (S1) of 2023 July 13, 06:40 UT 4 minutes, and for source 2 (S2) of 2023 July 13, 07:17 UT 6 minutes (see solid lines in panels (b) and (c)). The release time of S1 coincides with the onset time of a type-III radio burst. The derived path lengths are for S1 and for S2, indicating distinctly different magnetic connectivity for the two sources. For event 2, as shown in the right-hand panel of Fig. 3, a similar analysis is performed. Onset time for the sunward telescope is determined using the Poisson-CUSUM method, while for the anti-sunward telescope it is estimated visually (by averaging the fit results from ten independent visual examinations), due to the high background intensity resulting from the first injection in this event. The VDA results show injection times of 2024 July 22, 23:45 UT 6 minutes for source 1 and 2024 July 23, 00:15 UT 8 minutes for source 2. The corresponding path lengths are for S1 and for S2. Due to the strong scattering, the path length for this event might be less robust, especially for the S2. Based on these results, the derived onset times of S1 are reasonably consistent with the radio burst timing around 2024-07-22 23:45 UT and with around minutes delay for S2. For both events, the path lengths derived for sunward and anti-sunward particles are significantly greater than the nominal Parker spiral path length at for event 1 and at for event 2. By combining the estimated path lengths with the magnetic helicity of the identified flux rope, it is possible to infer the topology and spatial extent of the flux rope encountered in situ. An example of the calculation for event 1 is provided in Appendix B.
@@
The different release times indicate possible different origins for the source particles for both events. The close temporal coincidence between the release time determined from the sunward telescope and the onset of the type III radio burst suggests that the anti-sunward streaming particles might be promptly accelerated and released during flare-related magnetic reconnection. In contrast, a release time delayed by approximately minutes beyond the onset of the type III radio burst, accompanied by higher particle intensities as observed from the anti-sunward telescope, may indicate that particles originate from CME-driven shock acceleration. Numerous previous studies have shown that shock acceleration typically becomes efficient only when the CME-driven shock reaches high altitudes in the corona (e.g., Kahler, 1994; Cliver et al., 2004; Reames, 2009; Rouillard et al., 2011, 2012). This can be explained by two main processes. First, the formation of the shock itself: as the CME undergoes acceleration in the lower corona and exceeds the local fast magnetosonic speed, it drives the formation of a shock. Second, once the shock forms, particle acceleration generally proceeds via diffusive shock acceleration, where the acceleration time of high-energy particles depends on shock properties and wave intensity near the shock. Both factors likely contribute to the late release of high-energy particles. Several studies have investigated the shock formation heights and release heights of high-energy SEPs to better constrain the coronal acceleration environment. Gopalswamy et al. (2013) determined the shock formation heights at the onset time of metric Type II bursts for 32 SEP events between 2010 and 2012, finding shock heights ranging from to with a mean value of . Kouloumvakos et al. (2015) analyzed 65 SEP events and estimated the proton release heights to range from to , with an average of . More recently, Ameri et al. (2024) examined 58 high-energy proton events and found that events with typically exhibited proton release heights above , suggesting that less energetic SEP events are more likely associated with acceleration processes occurring higher in the corona. Therefore, the delayed release time of sunward particles found in this work likely reflects the time required for the CME-driven shock to develop and become sufficiently strong to accelerate high-energy protons and release them onto magnetic field lines.
@@
In applying the traditional VDA, obtaining the onset time is often the most challenging step. For our events, we employ both the Poisson-CUSUM method and visual inspection to estimate the onset times, which inevitably introduces some degree of uncertainty. Nevertheless the solid black line and purple lines in Figure 3 (panels (b) and (c)) provide reasonable proxies for the onset times. Another issue with VDA is the assumption that particles of different energies are released from the source at the same time. Studies of energetic electrons in impulsive events suggest that such an assumption is reasonable with an uncertainty of minutes (Zhao et al., 2019; Wu et al., 2023). For our events, we consider ions rather than electrons. For both events, if the first episode (S1) is due to flare acceleration, we expect the uncertainty obtained in the release time to be around a few minutes. For the second episode (S2), if the acceleration process is due to shock acceleration, then compared to that at flares, the acceleration could take more time. This effect is even more pronounced for higher-energy ions. The net effect is that the release time obtained will have an uncertainty in the order of the acceleration time scale of the maximum proton energy used in the VDA analysis. The maximum energy achievable at the shock front is determined by equating the particle acceleration time scale with the shock dynamic time scale (e.g., Li, 2017), where is the particle diffusion coefficient and is the upstream plasma speed in the shock frame. As the shock evolves quickly near the Sun, the shock dynamic time scale quickly drops, so is the maximum particle energy (Li et al., 2003; Li, 2017). The maximum particle energy considered in both events is approximately MeV. While direct measurements of the shock formation height and its evolution are not clear for the two events, we adopt representative values based on previous statistical studies as discussed above. We assume the shock forms at and propagates outward to a release height of , maintaining relatively stable shock characteristics (compression ratio, shock speed, and turbulence level upstream of the shock) during this interval. Using an average shock speed of for both events, we estimate the available acceleration time to be approximately minutes. Since the fitted release time of the second episodes trail those of the first episodes in both events by minutes, it is therefore plausible that both episodes correspond to the same “eruption” process near the Sun, with the first episodes having shorter acceleration time, consisting with flare acceleration; and the second episodes having longer acceleration time, consisting with shock acceleration. To support this, we consider the spectrum of both episodes below as the flare acceleration and shock acceleration likely yield different spectral indices.
@@
To further characterize the potential differences between source 1 and source 2 in both events, Fig. 4 presents the time-averaged proton energy spectra from the sunward and anti-sunward telescopes, calculated within a one-hour window following the onset times indicated by the dashed black and magenta boxes for S1 and S2 in Fig. 3. The duration of one-hour is chosen to increase statistical significance and to avoid overlap between the windows of S1 and S2. We note that shortening the time window does not significantly change the spectral features. For clarity, the spectra from the north and south telescopes are omitted. We focus only on proton energies above with clearly different anisotropies, because the low-energy particle intensity observed in the anti-sunward telescope may be contaminated by a pre-existing anti-sunward streaming particle population. The spectral index at the high-energy end is often used as an indicator of the acceleration efficiency, although it can also be influenced by transport and escape processes (e.g., Mewaldt et al., 2012; Desai et al., 2016).
@@
In event 1, as shown in the left-hand panel of Fig. 4, the proton intensity from S1 (black) is predominantly observed by the sunward telescope, and exhibits a power-law distribution of . Although the anti-sunward telescope detects only a very low intensity, it still exceeds the pre-event background level. For S2 (magenta), however, the particle intensity is higher in the anti-sunward telescope across all energy channels. Notably, the spectral index of source 2 in the anti-sunward telescope is harder (approximately ) than that of source 1 in the sunward telescope (approximately ), which is close to the limit of theoretical spectral index for a high compression ratio () in diffusive shock acceleration theory. CME-driven shock acceleration is generally more efficient at producing higher SEP intensity compared to flare acceleration (e.g., Li & Zank, 2005; Reames, 2013). The significantly higher particle intensity and harder spectral index both further support that the particles from S2 are likely to have been accelerated by the CME-driven shock.
@@
In contrast, in the right-hand panel of Fig. 4, event 2 exhibits much stronger scattering, as evidenced by the detection of particles in all telescopes in Fig. 3. This indicates a more isotropic distribution compared to event 1 and results in nearly identical spectral indices for the sunward and anti-sunward telescopes. Nevertheless, distinctions between the two sources remain evident. For S1, the particle intensity is approximately two orders of magnitude lower than that of S2, and the spectral index is softer, around . During the first injection (S1), the intensity observed by the sunward telescope is roughly twice that of the anti-sunward telescope. Conversely, during the second injection (S2), the intensity at the anti-sunward telescope is about twice that at the sunward telescope. These comparisons also support the conclusion that particles from the second injection predominantly stream toward the Sun and originate from the shock.
@@
Figure 5 presents a two-dimensional schematic illustration of the possible origin of bidirectional SEPs within a magnetic flux rope, as discussed in this study. In this scenario, the observer is situated inside a flux rope rooted at the Sun, with both legs anchored in the lower corona and one leg connecting directly to the flare site. Following the eruption, anti-sunward SEPs originating from flare acceleration reach the observer first via a shorter path length, while sunward SEPs, associated with stronger CME-driven shock acceleration, arrive later after traveling a longer path. Particles accelerated at the shock front can also propagate along field lines connecting to the flare site later on. However, to observe a distinct bidirectional velocity dispersion, the particle intensity along the field line connecting to the sunward telescope should be lower than that connecting to the anti-sunward telescope. This difference may reflect the distinct shock geometries encountered by each leg. For example, quasi-parallel shocks may have higher injection efficiency of seed particles compared to quasi-perpendicular shocks (Zank et al., 2006; Li et al., 2012; Ding et al., 2023). A detailed assessment of the shock geometry of these two events would require comprehensive shock modeling (e.g, Li et al., 2021; Ding et al., 2022, 2025), which is beyond the scope of this study but could be addressed in future work. We note that the observer could also be located at the opposite leg of the flux rope, with this leg connected to the flare site. Due to the lack of multi-spacecraft observations, we do not further discuss the possible orientation or expansion direction of the flux rope. This configuration demonstrates how a flux rope enables the simultaneous observation of SEPs from different sources, characterized by distinct release times and path lengths, as discussed in our results.

### v2

@@ 3 Results
Compared to event 1, event 2 exhibits more complex features, as shown in the right-hand panels of Fig. 2. There are data gaps in the SWA measurements during the pre-event period on July 22, 2024. The proton dynamic spectra in panel (a) also display two distinct tracks of velocity dispersion. Bidirectional streaming characteristics are again apparent in the first-order anisotropy shown in panel (b). In this event, the magnetic field polarity is predominantly inwards, so negative first-order anisotropy indicates particles propagating away from the Sun, whereas later positive values correspond to sunward streaming. This interpretation is reflected in panel (c), where the highest intensity of protons initially appears at pitch angles around (as measured by the sunward telescope) but later shifts to around (anti-sunward telescope). Panels (c) and (d) reveal that pitch angle distribution in this event exhibits more pronounced fluctuations than in event 1. The pitch angle distribution of protons in panel (d) shows a feature that is similar to event 1, during the period of 2:00 to 4:00 UT, the highest intensity of protons is measured with pitch angle around but the highest intensity of protons appears at pitch angles around , indicating the bidirectional SEP propagation. Unlike event 1, the bidirectional electron feature measured by EAS is less evident in the pitch angle distribution shown in panel (e). Although the phase space densities near both and are elevated compared to , the peak at is more pronounced in the inward magnetic field configuration. The magnetic field and solar wind data in panel (f)-(i) reveal substantial fluctuations, making it challenging to identify SFRs directly using PyGS. Only a few short-duration and unrelated SFRs are detected. Notably, analysis of the normalized reduced magnetic helicity reveals the presence of a highly helical magnetic structure with a characteristic scale of approximately minutes, spanning from around 00:00 to 06:00 UT on July 23, 2024. Although the precise boundaries are difficult to determine using the wavelet method, the combined results of and in Appendix A.2 indicate that Solar Orbiter crossed a potentially Alfvénic SFR embedded within a turbulent solar wind environment near the onset of the eruption. Unlike event 1, where the SFR was encountered before the eruption, in event 2, the encounter with the SFR is unclear. It seems to occur during the onset phase, which could allow for the observation of bidirectional SEPs. Recent studies have emphasized the prevalence of such SFRs in the slow solar wind (e.g. Chen et al., 2021; Shi et al., 2021), where they exhibit a high degree of Alfvénicity. The possible Alfvénic flux ropes may have a significant impact on SEP transport by affecting both the propagation path and scattering conditions. Unfortunately, we did not find any relevant eruption information associated with this possible SFR in the DONKI database. If the small eruption occurred precisely on the far side of the Sun, it is possible that no evidence would be detectable in coronagraph observations near Earth.
@@
To further understand the transport processes in the two events, we employed VDA to determine the release times and path lengths of the two sources. In the left-hand panels of Fig. 3, the upper panel displays the type III radio burst observed by RPW during event 1. Only a single burst is evident during the onset phase, beginning at approximately 06:41 UT from the Sun, indicated by the vertical line. The lower four panels show proton dynamic spectra from the four HET and EPT telescopes, highlighting the features of velocity dispersion. Two distinct velocity dispersion tracks can be readily identified by visual inspection in both the sunward and anti-sunward telescopes. Notably, the earliest arriving particles detected by the sunward telescopes are scarcely observed by the anti-sunward telescope during the onset. In addition, only a small number of particles are detected in the north and south telescopes at pitch angles near , indicating very limited scattering during the transport process. Compared to the proton measurements, velocity dispersion features are less pronounced in the electron data, likely due to high-level background noise. Therefore, we present the electron measurements by EPT in the Appendix A.1 and do not perform VDA for electrons in this study.
@@
Onset times are determined independently for the sunward and anti-sunward telescopes using the Poisson-CUSUM method (Lucas, 1985). VDA yields an injection time for source 1 (S1) of July 13, 2023, 06:40 UT 4 minutes, and for source 2 (S2) of July 13, 2023, 07:17 UT 6 minutes (see the solid lines in panels b and c). The release time of S1 coincides with the onset time of a type III radio burst. The derived path lengths are for S1 and for S2, indicating distinctly different magnetic connectivity for the two sources. For event 2, as shown in the right-hand panel of Fig. 3, a similar analysis is performed. Onset time for the sunward telescope is determined using the Poisson-CUSUM method, while for the anti-sunward telescope it is estimated visually (by averaging the fit results from ten independent visual examinations), due to the high background intensity resulting from the first injection in this event. The VDA results show injection times of July 22, 2024, 23:45 UT 6 minutes for source 1 and July 23, 2024, 00:15 UT 8 minutes for source 2. The corresponding path lengths are for S1 and for S2. Due to the strong scattering, the path length for this event might be less robust, especially for the S2. Based on these results, the derived onset times of S1 are reasonably consistent with the radio burst timing on July 22, 2024, at around 23:45 UT and with around minutes delay for S2. For both events, the path lengths derived for sunward and anti-sunward particles are significantly greater than the nominal Parker spiral path length at for event 1 and at for event 2. By combining the estimated path lengths with the magnetic helicity of the identified flux rope, it is possible to infer the topology and spatial extent of the flux rope encountered in situ. An example of the calculation for event 1 is provided in Appendix B.
@@
The different release times indicate possible different origins for the source particles for both events. The close temporal coincidence between the release time determined from the sunward telescope and the onset of the type III radio burst suggests that the anti-sunward streaming particles might be promptly accelerated and released during flare-related magnetic reconnection. In contrast, a release time delayed by approximately minutes beyond the onset of the type III radio burst, accompanied by higher particle intensities as observed from the anti-sunward telescope, may indicate that particles originate from CME-driven shock acceleration. Numerous previous studies have shown that shock acceleration typically becomes efficient only when the CME-driven shock reaches high altitudes in the corona (e.g. Kahler, 1994; Cliver et al., 2004; Reames, 2009; Rouillard et al., 2011, 2012). This can be explained by two main processes. First, the formation of the shock itself: as the CME undergoes acceleration in the lower corona and exceeds the local fast magnetosonic speed, it drives the formation of a shock. Second, once the shock forms, particle acceleration generally proceeds via diffusive shock acceleration, where the acceleration time of high-energy particles depends on shock properties and wave intensity near the shock. Both factors likely contribute to the late release of high-energy particles. Several studies have investigated the shock formation heights and release heights of high-energy SEPs to better constrain the coronal acceleration environment. Gopalswamy et al. (2013) determined the shock formation heights at the onset time of metric type II bursts for 32 SEP events between 2010 and 2012, finding shock heights ranging from to with a mean value of . Kouloumvakos et al. (2015) analysed 65 SEP events and estimated the proton release heights to range from to , with an average of . More recently, Ameri et al. (2024) examined 58 high-energy proton events and found that events with typically exhibited proton release heights above , suggesting that less energetic SEP events are more likely associated with acceleration processes occurring higher in the corona. Therefore, the delayed release time of sunward particles found in this work likely reflects the time required for the CME-driven shock to develop and become sufficiently strong to accelerate high-energy protons and release them onto magnetic field lines.
@@
In applying the traditional VDA, obtaining the onset time is often the most challenging step. For our events, we employed both the Poisson-CUSUM method and visual inspection to estimate the onset times, which inevitably introduces some degree of uncertainty. Nevertheless, the solid black line and purple lines in Fig. 3 (panels b and c) provide reasonable proxies for the onset times. Another issue with VDA is the assumption that particles of different energies are released from the source at the same time. Studies of energetic electrons in impulsive events suggest that such an assumption is reasonable with an uncertainty of minutes (Zhao et al., 2019; Wu et al., 2023). For our events, we considered ions rather than electrons. For both events, if the first episode (S1) is due to flare acceleration, we expect the uncertainty obtained in the release time to be around a few minutes. For the second episode (S2), if the acceleration process is due to shock acceleration, then compared to that at flares, the acceleration could take more time. This effect is even more pronounced for higher-energy ions. The net effect is that the release time obtained will have an uncertainty in the order of the acceleration timescale () of the maximum proton energy used in the VDA analysis. The maximum energy achievable at the shock front is determined by equating the particle acceleration timescale () with the shock dynamic timescale (e.g. Li, 2017), where is the particle diffusion coefficient and is the upstream plasma speed in the shock frame. As the shock evolves quickly near the Sun, the shock dynamic timescale quickly drops, so is the maximum particle energy (Li et al., 2003; Li, 2017). The maximum particle energy considered in both events is approximately MeV. While direct measurements of the shock formation height and its evolution are not clear for the two events, we adopted representative values based on previous statistical studies as discussed above. We assumed the shock forms at and propagates outwards to a release height of , maintaining relatively stable shock characteristics (compression ratio, shock speed, and turbulence level upstream of the shock) during this interval. Using an average shock speed of for both events, we estimate the available acceleration time to be approximately minutes. Since the fitted release time of the second episodes trail those of the first episodes in both events by minutes, it is therefore plausible that the two episodes correspond to the same ‘eruption’ process near the Sun, with the first episodes having shorter acceleration time, consisting with flare acceleration; and the second episodes having longer acceleration time, consisting with shock acceleration. To support this, we considered the spectrum of both episodes (see below) as the flare acceleration and shock acceleration likely yield different spectral indices.
@@
To further characterize the potential differences between source 1 and source 2 in both events, Fig. 4 presents the time-averaged proton energy spectra from the sunward and anti-sunward telescopes, calculated within a one-hour window following the onset times indicated by the dashed black and magenta boxes for S1 and S2 in Fig. 3. The duration of one hour is chosen to increase statistical significance and to avoid overlap between the windows of S1 and S2. We note that shortening the time window does not significantly change the spectral features. For clarity, the spectra from the north and south telescopes are omitted. We focused only on proton energies above with clearly different anisotropies because the low-energy particle intensity observed in the anti-sunward telescope may be contaminated by a pre-existing anti-sunward streaming particle population. The spectral index at the high-energy end is often used as an indicator of the acceleration efficiency, although it can also be influenced by transport and escape processes (e.g. Mewaldt et al., 2012; Desai et al., 2016).
@@
In event 1, as shown in the left-hand panel of Fig. 4, the proton intensity from S1 (black) is predominantly observed by the sunward telescope, and exhibits a power-law distribution of . Although the anti-sunward telescope detects only a very low intensity, it still exceeds the pre-event background level. For S2 (magenta), however, the particle intensity is higher in the anti-sunward telescope across all energy channels. Notably, the spectral index of source 2 in the anti-sunward telescope is harder (approximately ) than that of source 1 in the sunward telescope (approximately ), which is close to the limit of theoretical spectral index for a high compression ratio () in diffusive shock acceleration theory. CME-driven shock acceleration is generally more efficient at producing higher SEP intensity compared to flare acceleration (e.g. Li & Zank, 2005; Reames, 2013). The significantly higher particle intensity and harder spectral index are further evidence that the particles from S2 were accelerated by the CME-driven shock.
@@
In contrast, in the right-hand panel of Fig. 4, event 2 exhibits much stronger scattering, as evidenced by the detection of particles in all telescopes in Fig. 3. This indicates a more isotropic distribution compared to event 1 and results in nearly identical spectral indices for the sunward and anti-sunward telescopes. Nevertheless, distinctions between the two sources remain evident. For S1, the particle intensity is approximately two orders of magnitude lower than that of S2, and the spectral index is softer, around . During the first injection (S1), the intensity observed by the sunward telescope is roughly twice that of the anti-sunward telescope. Conversely, during the second injection (S2), the intensity at the anti-sunward telescope is about twice that at the sunward telescope. These comparisons also support the conclusion that particles from the second injection predominantly stream towards the Sun and originate from the shock.
@@
Figure 5 presents a two-dimensional schematic illustration of the possible origin of bidirectional SEPs within a magnetic flux rope, as discussed in this study. In this scenario, the observer is situated inside a flux rope rooted at the Sun, with both legs anchored in the lower corona and one leg connecting directly to the flare site. Following the eruption, anti-sunward SEPs originating from flare acceleration reach the observer first via a shorter path length, while sunward SEPs, associated with stronger CME-driven shock acceleration, arrive later after travelling a longer path. Particles accelerated at the shock front can also propagate along field lines connecting to the flare site later on. However, to observe a distinct bidirectional velocity dispersion, the particle intensity along the field line connecting to the sunward telescope should be lower than that connecting to the anti-sunward telescope. This difference may reflect the distinct shock geometries encountered by each leg. For example, quasi-parallel shocks may have a higher injection efficiency of seed particles compared to quasi-perpendicular shocks (Zank et al., 2006; Li et al., 2012; Ding et al., 2023). A detailed assessment of the shock geometry of these two events would require comprehensive shock modelling (e.g. Li et al., 2021; Ding et al., 2022, 2025), which is beyond the scope of this study but could be addressed in future work. We note that the observer could also be located at the opposite leg of the flux rope, with this leg connected to the flare site. Due to the lack of multi-spacecraft observations, we do not further discuss the possible orientation or expansion direction of the flux rope. This configuration demonstrates how a flux rope enables the simultaneous observation of SEPs from different sources, characterized by distinct release times and path lengths, as discussed in our results.